OpenStudy (anonymous):

Solve for x: 2 over 3 (x - 2) = 4x. negative 2 over 5 negative 5 over 2 2 over 5 5 @HomeschoolGrad @abb0t @Questionman @CashBoy27

OpenStudy (anonymous):

@SamuelAlden917

OpenStudy (anonymous):

@roro14 @shelbygt520

OpenStudy (anonymous):

when u say over do u mean fraction?

OpenStudy (anonymous):

Can you send the link to the page its confusing

OpenStudy (anonymous):

yea over meaning a fraction

OpenStudy (anonymous):

oh, okay. I gochu. 2/3 (x-2) = 4x 2/3x - 4/3 = 4x -4/3=10/3 x= -5/2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i beleive x= -5/2 is right so what @Decii_Jade said

OpenStudy (anonymous):

can someone from here help me with my biology

OpenStudy (anonymous):

your welcome and thanks @taffalousswagg_bru